Imagine that you are on interplanetary vessel coasting through a space to a rendezvous with Neptune. You discover that an immediate course correction is necessary. Otherwise the ship will not fall into orbit around Neptune, but will fly off into interstellar space! Your onboard computers indicate that the ship must gain 100.0m/s of forward speed to get back on course. Suppose the mass of the craft is exactly 2.0 x 10^4 kg. The rocket produce a constant force of 2.500 x 10^4 N.



A. When the rockets are fired, how fast will the ship accelerate?



B. How long will it take for the forward velocity of the ship to increase by the required 100.0m/s?



C. What will be the impulse that produces this course correction?

1
Expert's answer
2021-11-30T17:44:31-0500

a)


"a=\frac{F}{m}=\frac{2.5}{2}=1.25\frac{m}{s^2}"

b)


"t=\frac{100}{1.25}=80\ s"

c)


"I=25000(80)=2\cdot10^6\ Ns"